The perimeter of a rectangle is 220 inches. The ratio of its length to its width is 7:3. Find the area of the rectangle. I need the answer in square inches.
Answer by ikleyn(52781)

.

The length is 7x units long; the width is 3x units long.


The perimeter equation  2L + 2W = 220  takes the form


    2*(7x) + 2*(3x) = 220,

or

    14x + 6x = 220,

       20x   = 220

         x   = 220/20 = 11 inches.


Therefore, the length is  7*11 = 77 inches;  the width is  3*11 = 33 inches

and the area of the rectangle is  77*33 = 847  square inches.    ANSWER

Solved and thoroughly explained.